# PosIR: Position-Aware Heterogeneous Information Retrieval Benchmark

### v1

@@ 3 The PosIR Benchmark
Table 1: Summary of PosIR dataset statistics.
Table 1 provides a statistical overview of PosIR, which comprises 310 retrieval datasets spanning 10 languages [^8] Arabic, Chinese, German, English, French, Italian, Korean, Portuguese, Russian, Spanish. and 31 domains [^9] The 31 domains are grouped into the following high-level categories: Technology and Science, Industry and Economics, Health and Medicine, Society and Law, Entertainment and Culture, Finance, and Miscellaneous.. All datasets are following the same format as BEIR, i.e. corpus, queries and qrels, which are all available in the Hugging Face Hub. [^10] https://huggingface.co/datasets/infgrad/PosIR-Benchmark-v1 More details on PosIR are available in Appendix A.
@@ 3.1 Positional Distribution Analysis
In the position-aware fine-grained qrels $\mathcal{R}_{+}^{\prime}=\{q_{j},d_{j}^{+},\mathrm{ref}_{j},\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})\}_{j=1}^{|\mathcal{Q}^{\prime}|}$, we explicitly incorporate positional information through $\mathrm{pos}_{j}$. To analyze the distributional characteristics of this positional information, we aggregate the qrels across all domains separately for the English and Chinese datasets. For each query-document pair, we compute the mean position of $\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})$ and normalize it by the character length of the corresponding positive document $d_{j}^{+}$, yielding a relative position in the range $[0,1]$. We then discretize the relative positions into 20 equal-width bins and count the number of queries falling into each bin. The resulting positional distributions are illustrated in Figure 2, confirming that reference spans are broadly distributed across document positions, without a pronounced bias toward early segments as observed in MS-MARCO Hofstätter et al. (2021b); Coelho et al. (2024). Such a broad distribution is crucial for evaluating retrieval models’ sensitivity to position bias, as it ensures that relevance signals occur throughout the document rather than being dominated by leading content.
@@ 3.2 Positioning of PosIR
We position PosIR as a multilingual, position-aware retrieval benchmark designed to expose position bias that is invisible to existing evaluation suites. Unlike popular retrieval benchmarks such as MTEB Muennighoff et al. (2023) and C-MTEB Xiao et al. (2024), which predominantly focus on short documents (typically under 512 tokens), PosIR explicitly incorporates documents with diverse and realistic length distributions, where relevant information may appear at arbitrary positions within the documents. A key distinguishing feature of PosIR is the fine-grained, position-aware relevance annotations. Rather than treating document relevance as a holistic property, PosIR associates each query-document pair with a localized reference span, allowing direct analysis of how retrieval models respond to the positional distribution of relevant information. This unique characteristic provides a solid empirical foundation for studying position bias in retrieval models, which is largely unexplored in existing benchmarks.
Model MMTEB PosIR Q1(512) Q2(1024) Q3(1536) Q4(2048) gte-multilingual-base Zhang et al. (2024) 57.16 47.37 61.28 48.79 39.21 32.01 bge-m3 Chen et al. (2024) 54.60 43.22 57.16 43.48 35.13 29.72 Qwen3-Embedding-0.6B Zhang et al. (2025) 64.65 53.63 62.93 54.41 48.11 43.32 inf-retriever-v1-1.5b Yang et al. (2025b) 62.96 58.81 67.82 58.40 53.90 51.20 Qwen3-Embedding-4B Zhang et al. (2025) 69.60 62.26 71.63 62.91 56.74 50.96 inf-retriever-v1 Yang et al. (2025b) 66.48 65.01 74.71 65.03 59.82 54.68 NV-Embed-v2 Lee et al. (2025) 56.72 45.02 70.48 47.12 26.33 16.27 llama-embed-nemotron-8b Babakhin et al. (2025) 68.69 64.09 75.76 64.16 57.47 52.28 Qwen3-Embedding-8B Zhang et al. (2025) 70.88 64.08 72.68 64.33 59.00 53.87 KaLM-Embedding-12B Zhao et al. (2025b) 75.66 51.87 74.01 54.64 35.11 27.65 Spearman Rank Correlation Coefficient (p-value) 0.62 (p=0.05) 0.73 (p=0.01) 0.71 ((p=0.02) 0.44 (p=0.2) 0.39 (p=0.2)
Table 2: nDCG@10 $\uparrow$ performance of 10 multilingual retrieval models on MMTEB (Retrieval Task) and PosIR. MMTEB scores are sourced from the official leaderboard. For PosIR, the results are first weighted-averaged across 31 domains and then macro-averaged across 10 languages. Q1–Q4 represent query buckets partitioned by the token length of positive documents (512-token intervals). Spearman rank correlation coefficients between MMTEB and PosIR (overall and per length bucket) are reported at the bottom. “KaLM-Embedding-12B” denotes KaLM-Embedding-Gemma3-12B-2511. The detailed results for each language in PosIR are available in Table 10.
@@ 4 Experiments
RQ3: What internal model behaviors and representational dynamics contribute to the emergence of position bias in neural retrieval models?
@@ 4.1 Comparison with MMTEB (RQ1)
We adopt MMTEB (Retrieval Task) as a representative multilingual retrieval benchmark. We evaluate 10 popular IR models on both MMTEB and PosIR using nDCG@10, and compute the Spearman rank correlation coefficient between model rankings on the two benchmarks as a measure of consistency. To further analyze the effect of document length, we partition queries in PosIR into four length buckets (Q1–Q4) with 512-token intervals, based on the token length of their positive documents $d_{j}^{+}$ in the qrels $\mathcal{R}_{+}^{\prime}$. For translated datasets, token lengths are measured based on their original English counterparts to ensure consistent bucket assignment across languages. We then compute Spearman correlations between MMTEB and the average performance of models within each length bucket.
@@
A distinct pattern is observed for NV-Embed-v2 and KaLM-Embedding-Gemma3-12B-2511, which achieve competitive performance on MMTEB and short-document queries (Q1), but suffer substantial degradation in Q3 and Q4. This behavior is consistent with their reported training configurations Lee et al. (2025); Zhao et al. (2025a), as both models are primarily trained with short-context inputs (e.g., up to 512 tokens), limiting their ability to effectively encode long documents. Furthermore, NV-Embed-v2, while strong in Q1, degrades severely in longer-document buckets in multilingual settings. This aligns with the fact that it is trained predominantly on English data, suggesting that long-context cross-lingual generalization poses a significant challenge when both document length and language shift are combined.
@@
Table 3: Position Sensitivity Index (PSI) $\downarrow$ of 10 multilingual retrieval models on PosIR. In both multilingual retrieval and cross-lingual retrieval (translated queries retrieving English documents) settings, the results are first weighted-averaged across 31 domains and then macro-averaged across 10 languages. Q1–Q4 represent query buckets partitioned by the token length of positive documents (512-token intervals). “KaLM-Embedding-12B” denotes the “KaLM-Embedding-Gemma3-12B-2511” model, while NV-Embed-v2 improves the mean* pooling method with a latent attention layer. The detailed results for each language in PosIR are available in Table 10.
@@
Figure 4: Mean nDCG@10 scores of 10 IR models across 20 relative position bins in the French-to-English cross-lingual setting of PosIR. “KaLM-Embedding-12B” denotes KaLM-Embedding-Gemma3-12B-2511.
@@ 4.2 Patterns of Position Bias (RQ2)
Following Zeng et al. (2025b), we adopt the Position Sensitivity Index (PSI) as an intuitive diagnostic metric for quantifying position bias from a worst-case perspective. We divide the queries into four length buckets (Q1–Q4) according to the strategy in Section 4.1. Within each bucket, we calculate the average nDCG@10 scores across the 20 equal-width relative position bins (denoted as $\mathbf{s}=\{s_{1},\ldots,s_{20}\}$), following the discretization described in Section 3.1. Formally, PSI is defined as $\mathrm{PSI}=1-\frac{\min(\mathbf{s})}{\max(\mathbf{s})}$. A lower PSI indicates that the model’s performance is more consistent across document positions, reflecting reduced sensitivity to the location of relevant content. For a more detailed discussion of the PSI metric, please refer to the original work.
@@
As shown in Table 3, we observe a clear trend: position bias tends to increase with document length across both multilingual and cross-lingual retrieval tasks. [^11] We consider a cross-lingual retrieval setting where non-English queries are used to retrieve an English corpus, which is a common scenario. For short documents (Q1), most models exhibit relatively low PSI values, suggesting that position bias is minimal when processing shorter inputs. However, as document length increases (Q3 and Q4), there is a marked rise in PSI for several models, indicating heightened sensitivity to position in long-context retrieval scenarios. Moreover, cross-lingual retrieval exhibits a lower overall PSI compared to the monolingual setting. We hypothesize this is partly because the generally lower performance in cross-lingual tasks compresses the score range, thereby reducing the discriminative power of the PSI metric across position bins. We also observe that for certain models, the PSI in Q3 or Q4 is counter-intuitively lower than in Q2. We attribute this to the models’ failure to effectively encode longer document representations, leading to uniformly degraded performance where position bias is no longer the dominant factor in score fluctuations. Moreover, we observe no clear correlation between positional sensitivity and architectural factors such as model size, vector dimension, attention mechanism (bidirectional / causal), or pooling strategy (CLS / last / mean). [^12] However, the latent attention layer preceding the mean pooling strategy may make NV-Embed-v2 more unique. This suggests that these architectural choices alone may not be the primary determinants of position bias, or their effects are overshadowed by training data distributions.
We visualize the fine-grained model performance across 20 relative position bins in Figures 3 and 4. Additional figures are available in the official GitHub repository. [^13] https://github.com/Ziyang1060/PosIR/tree/main/figs As shown in these figures, most models exhibit a generic pattern of primacy bias, where retrieval performance degrades as the relevant information moves towards the end of the document. An unexpected case is NV-Embed-v2, which exhibits a recency bias skewed towards the end of the document, diverging from the typical patterns observed in other models.
4.3 Mechanisms of Position Bias (RQ3)
@@
To better understand the mechanisms underlying the observed position bias, we conduct exploratory analyses on the internal behaviors of two representative models: Qwen3-Embedding-8B (exhibiting primacy bias) and NV-Embed-v2 (exhibiting recency bias). To quantitatively assess the contribution of tokens at different positions, we employ gradient-based saliency analysis Simonyan et al. (2014). Given a query $q$ and a relevant long document $d^{+}$ (with length $\geq$ 1024 tokens), where the reference span is located near the middle of the document (i.e., relative position within $[0.4,0.6]$), we compute the gradient of the cosine similarity score $s(q,d^{+})$ with respect to the input document embeddings. The magnitude of the gradient serves as a proxy for token importance, indicating how strongly each token influences the final document representation used for relevance matching. Specifically, for each document token $w_{i}$ at absolute position $i$, we compute the L2 norm of the gradient vector $\nabla_{w_{i}}s(q,d^{+})||_{2}$, which measures the sensitivity of the relevance score to that token. [^14] Special tokens at the sequence boundaries are excluded. To facilitate comparison across documents of varying lengths, we apply max normalization over all document tokens and rescale token positions to the range $[0,1]$ using linear interpolation, discretized into 100 relative position bins. We aggregate the saliency over 3,131 English query-document pairs that satisfy the length and position constraints to ensure statistical robustness.
Figure 5: Gradient-based saliency maps comparing the internal attention dynamics of Qwen3-Embedding-8B and NV-Embed-v2. The x-axis represents the normalized relative position within a document, and the y-axis shows the normalized L2 norm of the gradients. Shaded regions indicate one standard deviation.
@@
We visualize the gradient-based saliency maps in Figure 5. The results reveal starkly contrasting internal behaviors between the two models. The saliency curve for Qwen3-Embedding-8B (Orange) is characterized by an extreme peak at the very beginning. This confirms that despite being a causal model, Qwen3 heavily relies on the initial tokens to establish the semantic context for the entire sequence. The subsequent sharp decay reflects the model’s inability to effectively propagate gradients from the middle sections, providing a mechanistic explanation for its tendency to ignore mid-document information. We hypothesize that the spike at the final position arises as a structural artifact of last-token pooling. In stark contrast, NV-Embed-v2 (Blue) exhibits a suppressed sensitivity to the beginning of the document, with gradients hovering near zero for the first 20% of tokens. Instead, it demonstrates a continuous rising trend starting from the mid-point (0.5) and peaking at the end. This J-shaped curve suggests that NV-Embed’s encoding mechanism prioritizes recent information, progressively overwriting or diluting early context. This explains its atypical recency bias and its failure to retrieve information located at the start of long documents.
5 Related Work
5.1 Position Bias in IR
Prior work has widely documented the existence of position bias in neural IR systems. Hofstätter et al. (2021b) identified that MS MARCO, a foundational IR dataset, is heavily skewed toward head-located relevant spans, leading fine-tuned models to inherit and amplify this position bias Jiang et al. (2021), which compromises the reliability of existing evaluations Rau et al. (2024). Recent studies confirm that such bias persists even in state-of-the-art embedding models Coelho et al. (2024); Fayyaz et al. (2025); Zeng et al. (2025b). Specifically, Zeng et al. (2025b) utilize answer start positions in SQuAD Rajpurkar et al. (2018) for bias analysis; however, SQuAD is constrained to English monolingual data and brief passages (avg. 117 words), hindering the generalization of findings to multilingual or long-context scenarios. Despite these insights, a standardized framework to rigorously evaluate position bias, especially within multilingual and long contexts, remains absent. Consequently, PosIR is proposed to bridge this gap.
5.2 Synthetic Data for IR
Leveraging LLMs for synthetic data generation has recently attracted growing attention in the IR community. Existing works can be broadly categorized into two primary use cases: training and evaluation. From a training perspective, LLMs are employed to address the scarcity of domain-specific or task-specific data. Methods like InPars Bonifacio et al. (2022) and Promptagator (Dai et al., 2023) prompt LLMs to generate synthetic queries for specific documents, facilitating effective retriever training in zero-shot settings. This paradigm has been further extended by Thakur et al. (2024); Wang et al. (2024) to support multilingual retrieval tasks. Regarding evaluation, recent studies have validated the reliability of synthetic test collections for benchmarking Rahmani et al. (2024). Khramtsova et al. (2024) leverage LLM-generated pseudo-labels and queries for unsupervised model selection on a target corpus. Furthermore, frameworks like AIR-Bench (Chen et al., 2025) provide automated pipelines to generate evaluation data for emerging domains efficiently. Inspired by these advancements, we extend the application of synthetic data generation to the specific challenge of position bias. Unlike prior works that focus on document-level relevance, we utilize LLMs to generate fine-grained, position-aware relevance annotations, which are crucial for isolating position bias. By incorporating rigorous quality control mechanisms, PosIR ensures the reliability of synthetic signals, enabling efficient and effective diagnosis of position bias in retrieval models.

### v2

@@ 4 The PosIR Benchmark
Table 2 provides a statistical overview of PosIR, which comprises 310 retrieval datasets spanning 10 languages [^6] Arabic, Chinese, German, English, French, Italian, Korean, Portuguese, Russian, Spanish. and 31 domains [^7] The 31 domains are grouped into the following high-level categories: Technology and Science, Industry and Economics, Health and Medicine, Society and Law, Entertainment and Culture, Finance, and Miscellaneous.. More details on PosIR are available in Appendix 0.A.
@@ 4.1 Positional Distribution Analysis
In the position-aware fine-grained qrels $\mathcal{R}_{+}^{\prime}=\{q_{j},d_{j}^{+},\mathrm{ref}_{j},\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})\}_{j=1}^{|\mathcal{Q}^{\prime}|}$, we explicitly incorporate positional information through $\mathrm{pos}_{j}$. To analyze the distributional characteristics of this positional information, we aggregate the qrels across all domains separately for the English and Chinese datasets. For each query-document pair, we compute the mean position of $\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})$ and normalize it by the character length of the corresponding positive document $d_{j}^{+}$, yielding a relative position in the range $[0,1]$. We then discretize the relative positions into 20 equal-width bins and count the number of queries falling into each bin. The resulting positional distributions are illustrated in Figure 2, confirming that reference spans are broadly distributed across document positions, without a pronounced bias toward early segments as observed in MS-MARCO [11, 5]. Such a broad distribution is crucial for evaluating retrieval models’ sensitivity to position bias, as it ensures that relevance signals occur throughout the document rather than being dominated by leading content. This also serves as evidence that our automated generation pipeline does not introduce systematic positional artifacts: if the pipeline itself were biased toward particular positions [16], the distribution in Figure 2 would exhibit a pronounced skew, which is not observed.
@@ 4.2 Positioning of PosIR
Existing retrieval benchmarks serve complementary purposes: MMTEB [18, 8] and BEIR [26] evaluate general-purpose embedding quality, MIRACL [36] targets multilingual retrieval, and AIR-Bench [4] meets the demand for efficient evaluation in emerging domains. However, all of them assign position-agnostic relevance labels, leaving position bias undiagnosed. PosIR fills this gap as a multilingual, position-aware retrieval benchmark. By associating each query with a localized reference span within length-diverse documents, PosIR enables controlled analysis of how retrieval performance varies with evidence position—a dimension largely unexplored in existing benchmarks.
Table 2: nDCG@10 $\uparrow$ performance of 10 multilingual retrieval models on MMTEB (Retrieval Task) and PosIR. MMTEB scores are sourced from the official leaderboard. For PosIR, the results are first weighted-averaged across 31 domains and then macro-averaged across 10 languages. Q1–Q4 represent query buckets partitioned by the token length of positive documents (512-token intervals). Spearman rank correlation coefficients between MMTEB and PosIR (overall and per length bucket) are reported at the bottom. The detailed results for each language in PosIR are available in Table 10.
Model MMTEB PosIR Q1(512) Q2(1024) Q3(1536) Q4(2048) gte-multilingual-base [35] 57.16 47.37 61.28 48.79 39.21 32.01 bge-m3 [3] 54.60 43.22 57.16 43.48 35.13 29.72 Qwen3-Embedding-0.6B [37] 64.65 53.63 62.93 54.41 48.11 43.32 inf-retriever-v1-1.5b 62.96 58.81 67.82 58.40 53.90 51.20 Qwen3-Embedding-4B [37] 69.60 62.26 71.63 62.91 56.74 50.96 inf-retriever-v1 66.48 65.01 74.71 65.03 59.82 54.68 NV-Embed-v2 [15] 56.72 45.02 70.48 47.12 26.33 16.27 llama-embed-nemotron-8b [1] 68.69 64.09 75.76 64.16 57.47 52.28 Qwen3-Embedding-8B [37] 70.88 64.08 72.68 64.33 59.00 53.87 KaLM-Embedding-12B [38] 75.66 51.87 74.01 54.64 35.11 27.65 Spearman Corr. 0.62 0.73 0.71 0.44 0.39 (p-value) (0.05) (0.01) (0.02) (0.20) (0.20)
@@ 5 Experiments
RQ3: What internal model behaviors and representational dynamics contribute to the emergence of position bias in dense retrieval models?
@@ 5.1 Comparison with MMTEB (RQ1)
We adopt MMTEB (Retrieval Task) as a representative multilingual retrieval benchmark. We evaluate 10 popular embedding models on both MMTEB and PosIR using nDCG@10, and compute the Spearman rank correlation coefficient between model rankings on the two benchmarks as a measure of consistency. To further analyze the effect of document length, we partition queries in PosIR into four length buckets (Q1–Q4) with 512-token intervals, based on the token length of their positive documents $d_{j}^{+}$ in the qrels $\mathcal{R}_{+}^{\prime}$. For translated datasets, token lengths are measured based on their original English counterparts to ensure consistent bucket assignment across languages. We then compute Spearman correlations between MMTEB and the average performance of models within each length bucket.
@@
A pronounced pattern is observed for NV-Embed-v2 and KaLM-Embedding-12B, which achieve competitive performance on MMTEB and short-document queries (Q1), but suffer substantial degradation in Q3 and Q4. This behavior is consistent with their reported training configurations [15, 38], as both models are primarily trained with short-context inputs (e.g., up to 512 tokens), limiting their ability to effectively encode long documents. Furthermore, NV-Embed-v2, while strong in Q1, degrades severely in longer-document buckets in multilingual settings. This aligns with the fact that it is trained predominantly on English data, suggesting that long-context cross-lingual generalization poses a significant challenge when both document length and language shift are combined.
Table 3: Position Sensitivity Index (PSI) $\downarrow$ of 10 multilingual retrieval models on PosIR. In both multilingual retrieval and cross-lingual retrieval (translated queries retrieving English documents) settings, the results are first weighted-averaged across 31 domains and then macro-averaged across 10 languages. NV-Embed-v2 improves the mean* pooling method with a latent attention layer. The detailed results for each language in PosIR are available in Table 10.
@@
Figure 4: Mean nDCG@10 scores of 10 IR models across 20 relative position bins in the French-to-English cross-lingual setting of PosIR.
@@ 5.2 Patterns of Position Bias (RQ2)
We adopt the Position Sensitivity Index (PSI) introduced in [33] as an intuitive diagnostic metric for quantifying position bias from a worst-case perspective. We divide the queries into four length buckets (Q1–Q4) according to the strategy in Section 5.1. Within each bucket, we calculate the average nDCG@10 scores across the 20 equal-width relative position bins (denoted as $\mathbf{s}=\{s_{1},\ldots,s_{20}\}$), following the discretization described in Section 4.1. Formally, PSI is defined as: $\mathrm{PSI}=1-\frac{\min(\mathbf{s})}{\max(\mathbf{s})}$. A lower PSI indicates that the model’s performance is more consistent across document positions, reflecting reduced sensitivity to the location of relevant content. For a more detailed discussion of the PSI metric, please refer to the original work.
@@
As shown in Table 3, we observe a clear trend: position bias tends to increase with document length across both multilingual and cross-lingual retrieval tasks. [^8] We consider a cross-lingual retrieval setting where non-English queries are used to retrieve an English corpus, which is a common scenario. For short documents (Q1), most models exhibit relatively low PSI values, suggesting that position bias is minimal when processing short inputs. However, as document length increases (Q3 and Q4), PSI rises markedly for several models, indicating heightened sensitivity to position in longer-context retrieval scenarios. Moreover, cross-lingual retrieval exhibits a lower overall PSI compared to the monolingual setting. We hypothesize this is partly because the generally lower performance in cross-lingual tasks compresses the score range, thereby reducing the discriminative power of the PSI metric across position bins. We also observe that for certain models, the PSI in Q3 or Q4 is counter-intuitively lower than in Q2. We attribute this to the models’ failure to effectively encode longer document representations, leading to uniformly degraded performance where position bias is no longer the dominant factor in score fluctuations. Moreover, we observe no clear correlation between positional sensitivity and architectural factors such as model size, vector dimension, attention mechanism (bidirectional / causal), or pooling strategy (CLS / last / mean). [^9] However, the latent attention layer preceding the mean pooling strategy may make NV-Embed-v2 more unique. This suggests that these architectural choices alone may not be the primary determinants of position bias, or their effects are overshadowed by training data distributions.
We visualize the fine-grained model performance across 20 relative position bins in Figure 3 and 4. Additional figures are available in the official GitHub repository. [^10] https://github.com/Ziyang1060/PosIR/tree/main/figs As shown in these figures, most models exhibit a common pattern of primacy bias, where retrieval performance degrades as the relevant information moves towards the end of the document. An unexpected case is NV-Embed-v2, which exhibits a recency bias skewed towards the end of the document, diverging from the typical patterns observed in other models.
5.3 Exploring Mechanisms of Position Bias (RQ3)
We present a preliminary mechanistic analysis to generate hypotheses about the internal behaviors underlying the observed retrieval biases, rather than to provide definitive causal explanations.
@@ 5.3 Exploring Mechanisms of Position Bias (RQ3)
We conduct exploratory analyses on the internal behaviors of two representative models: Qwen3-Embedding-8B (exhibiting primacy bias) and NV-Embed-v2 (exhibiting recency bias). To quantitatively assess the contribution of tokens at different positions, we employ gradient-based saliency analysis [24]. Given a query $q$ and a relevant long document $d^{+}$ (with length $\geq$ 1024 tokens), where the reference span is located near the middle of the document (i.e., relative position within $[0.4,0.6]$), we compute the gradient of the cosine similarity score $s(q,d^{+})$ with respect to the input document embeddings. The magnitude of the gradient serves as a proxy for token importance, indicating how strongly each token influences the final document representation used for relevance matching. Specifically, for each document token $w_{i}$ at absolute position $i$, we compute the L2 norm of the gradient vector $\nabla_{w_{i}}s(q,d^{+})||_{2}$, which measures the sensitivity of the relevance score to that token. [^11] Special tokens at the sequence boundaries are excluded. To facilitate comparison across documents of varying lengths, we apply max normalization over all document tokens and rescale token positions to the range $[0,1]$ using linear interpolation, discretized into 100 relative position bins. We aggregate the saliency over 3,131 English query-document pairs that satisfy the length and position constraints to ensure statistical robustness.
Figure 5: Gradient-based saliency maps comparing the token-level sensitivity of Qwen3-Embedding-8B and NV-Embed-v2. The x-axis represents the normalized relative position within a document, and the y-axis shows the normalized L2 norm of the gradients. Shaded regions indicate one standard deviation.
@@
We visualize the gradient-based saliency maps in Figure 5. The results reveal starkly contrasting internal behaviors between the two models. The saliency curve for Qwen3-Embedding-8B (Orange) is characterized by an extreme peak at the very beginning. The subsequent sharp decay suggests the model may struggle to effectively propagate gradient signal from later sections, offering a candidate mechanistic explanation for its tendency to under-weight mid-document information. We further conjecture that the spike at the final position arises as a structural artifact of last-token pooling, though causal verification remains to be established. In stark contrast, NV-Embed-v2 (Blue) exhibits suppressed sensitivity to the beginning of the document, with gradients remaining relatively low for the first 20% of tokens. Instead, it demonstrates a continuous rising trend starting from the mid-point (0.5) and peaking at the end. This J-shaped profile is consistent with the hypothesis that NV-Embed-v2’s encoding mechanism progressively prioritizes recent information, potentially overwriting or diluting early context. We conjecture this pattern underlies its atypical recency bias and its failure to retrieve information located at the start of long documents. We emphasize that the above analyses are correlational. Intervention experiments such as token ablation and positional perturbation would be needed to establish causality, which we leave to future work.